# Continuum-Field-Theoretical Construction of Conserved Currents, Non-Invertible Symmetries, and Mixed Anomalies in (3+1)D Non-Abelian Topological Order

### v2

@@ II Preliminaries of conserved currents and symmetry operators
In field theory, a symmetry is indicated by a conserved current, which in turn gives rise to a topological operator, namely the symmetry operator associated with the symmetry. From this perspective, a symmetry refers to the mathematical structure formed by these topological operators. In our discussion of generalized symmetries, we adopt the viewpoint that a higher-form symmetry is generated by topological operators originating from higher-form conserved currents. Moreover, this principle naturally extends to the construction of non-invertible symmetries, whose symmetry operators can be traced back to a special class of conserved currents, as will be discussed in Sec. III. Two representative examples of generalized symmetries in lattice systems are collected in Appendix A.
@@
where the real parameter $\alpha$ takes value in $0,2\pi)$ as a compact phase angle. The compactness of $\alpha$ arises from the quantization of $Q$. The existence of this global symmetry can thus be traced back to the presence of a conserved current. The conservation law ensures that $U_{\alpha}$ is invariant under smooth deformations of $\Sigma_{3}$, hence the terminology topological operator of the 0-form symmetry (i.e., global symmetry) that act on the entire quantum system.
The concept of higher-form symmetry, e.g., $p$-form symmetry, generalizes the above structure by allowing the conserved current to be a $(p+1)$-form. Correspondingly, the associated symmetry operator is supported on a $\left(D-1-p\right)$-dimensional submanifold in (D1+1)(D-1+1)D spacetime. As an illustrative example of higher-form symmetries, we consider the (3+1)(3+1)D Maxwell’s electromagnetic theory with action $S=\frac{1}{4}\int d^{4}xF_{\mu\nu}F^{\mu\nu}$, where $F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$ and $e$ is the gauge coupling constant; also see the similar discussion in, e.g., Refs [8, 1, 3, 6]. First, there exists a conserved $2$-form current satisfying $d\left(*J_{m}\right)=0$, with $J_{m}^{\mu\nu}=\frac{1}{4\pi}\epsilon^{\mu\nu\lambda\sigma}F_{\lambda\sigma}=\frac{1}{2\pi}\left(d\widetilde{A}\right)^{\mu\nu}$. Here $\widetilde{A}$ is defined by $d\widetilde{A}=*dA$. This continuity equation is equivalent to $dF=0$, namely the Bianchi identity. The conserved current $J_{m}$ gives rise to a “magnetic” $1$-form symmetry with symmetry operator
@@
Another illustrative example is provided by the mutual Chern–Simons theory, which serves as the effective field theory of the 22d topological order,
@@
The lattice realization of this field theory is the celebrated toric code model in 2d square lattice. Here $a^{1}$ and $a^{2}$ are $1$-form gauge fields. The equations of motion are $\frac{1}{\pi}da^{1}=0$ and $\frac{1}{\pi}da^{2}=0$. In analogy with the previous examples, we interpret these equations as continuity equations and introduce the conserved quantities $Q_{1}=\int_{\gamma}a^{1}$ and $Q_{2}=\int_{\gamma}a^{2}$. These quantities measure the number of topological excitations enclosed by the area bounded by a loop $\gamma$. They generate two minimal symmetry operators,
@@ III Field-theoretical construction of symmetry operators
In this section, we construct symmetry operators in effective field theories of 33d topological orders. We begin with the $\left(3+1\right)$D $\mathbb{Z}_{N}$ BFBF theory as a warm-up, which describes $\mathbb{Z}_{N}$ topological orders. We then turn to the main goal of this paper: identifying symmetry operators in the effective field theory of a class of 33d non-Abelian topological orders, namely Borromean-Rings (BR) topological orders. In particular, we show how conserved currents can be exploited to construct non-invertible symmetry operators.
A system with BR topological order [48, 49, 76, 79] exhibits nontrivial BR braiding statistics. A BR braiding process involves one particle and two loop excitations: the particle is moved around the two loops such that its spatial trajectory, together with the two loops, forms Borromean rings. Borromean rings are three closed curves in 33d space that are not pairwise linked, yet cannot be separated as a whole. Consequently, there is no net particle–loop braiding phase, since the particle trajectory is not linked with either loop individually. BR braiding is also distinct from multi-loop braiding because it necessarily involves a particle excitation. The corresponding TQFT is a $BF$ theory with an aabaab-type twisted term; see Eq. (14) below. Remarkably, BR topological order supports non-Abelian particle and loop excitations even when all excitations carry gauge charges and fluxes of an Abelian gauge group, making it an ideal playground for exploring generalized symmetries in $\left(3+1\right)$D systems.
In the following, we first derive the equations of motion (EoM) of gauge fields, from which we identify conserved currents. These conserved currents imply symmetries and allow us to construct the corresponding symmetry operators. We find two classes of conserved currents denoted as type-I and type-II respectively, as shown in Sec. III.3 and Sec. III.4. Type-I currents lead to group-like (i.e., invertible) symmetries. Type-II currents are conserved only under certain constraints such that corresponding conserved quantities are legitimate. Accordingly, the symmetry operators generated by these conserved quantities must be accompanied by projectors that enforce the constraints. Since projectors have no inverses, these symmetry operators are naturally non-invertible.
@@ III.1 Higher-form symmetries in $\left(3+1\right)$D $\mathbb{Z}_{N}$ BFBF theory
The topological field theories describing the 22d and 33d toric code models discussed above are $\mathbb{Z}_{2}$ BFBF theories. It is straightforward to generalize the gauge group from $\mathbb{Z}_{2}$ to a general $\mathbb{Z}_{N}$. The $\left(3+1\right)$D $\mathbb{Z}_{N}$ BFBF theory is described by
@@
where $J_{e}$ and $J_{m}$ are $3$-form and $2$-form conserved currents, respectively. Analogous to ordinary 0-form symmetries, these conserved currents give rise to conserved quantities
@@
We now explain the physical meaning of these conserved quantities. Consider $Q_{e}\left(\gamma\right)=\int_{\gamma}a$ as an example. From the correspondence between gauge theory and topological order, the operator $e^{{\rm i}\int_{\gamma}a}$ is the Wilson operator associated with a particle excitation carrying one unit of gauge charge, and $\gamma$ can be interpreted as the worldline of this particle. Acting with $e^{{\rm i}\int_{\gamma}a}$ on a state amounts to transporting the particle along $\gamma$ in spacetime. This process probes the gauge flux carried by loop excitations piercing the surface bounded by $\gamma$.
From the perspective of conserved currents, since $a=a_{\mu}dx^{\mu}$, the current takes the form $J_{e}=-\frac{1}{3!}\epsilon^{\mu\nu\lambda\rho}a_{\mu}dx^{\nu}\wedge dx^{\lambda}\wedge dx^{\rho}$. For comparison, in a $\left(3+1\right)$D system with a 0-form $U\left(1\right)$ symmetry, the particle current $j=j_{\mu}dx^{\mu}$ is a $1$-form current, whose Hodge dual is $j=j_{0}dx^{1}\wedge dx^{2}\wedge dx^{3}-j_{1}dx^{0}\wedge dx^{2}\wedge dx^{3}+j_{2}dx^{0}\wedge dx^{1}\wedge dx^{3}-j_{3}dx^{0}\wedge dx^{1}\wedge dx^{2}$. Taking $M_{3}$ to be the entire 33d spatial manifold, $\int_{M_{3}}*j$ integrates the particle-number density $j_{0}$ over $M_{3}$, yielding the total particle number. By direct analogy, $\int_{\gamma}*J_{e}$ can be interpreted as a generalized “symmetry charge” associated with the submanifold $\gamma$.
Recall that $\int_{\gamma}*J_{e}$ counts the gauge fluxes encircled by $\gamma$; in $\left(3+1\right)$D, this quantity is given by the linking number between $\gamma$ and the worldsheets $\sigma$ of loop excitations. Equivalently, it can be computed by counting the intersections between $\gamma$ and a three-dimensional volume bounded by $\sigma$ in spacetime. In this sense, $J_{e}$ describes the flow of $2$-dimensional objects in $\left(3+1\right)$D spacetime, in close analogy with $j$ describing the flow of 0-dimensional particles. These 22d objects are precisely the worldsheets of loop excitations. The physical interpretation of $Q_{m}\left(\sigma\right)$ follows in a completely analogous manner.
@@
The equations of motion imply that $\oint a\in\frac{2\pi}{N}\mathbb{Z}_{N}$, which in turn gives $NQ_{e}\left(\gamma\right)\in 2\pi\mathbb{Z}$, meaning that $q_{e}$ is equivalent to $q_{e}+N$. Furthermore, large gauge invariance under $\int_{\gamma}a\rightarrow\int_{\gamma}a+2\pi$ requires $q_{e}$ to be an integer. Consequently, $q_{e}$ takes values in $\mathbb{Z}/N\mathbb{Z}$. The symmetry operators satisfy
@@ III.2 Continuum topological field theory of Borromean-Rings topological order
To identify the conserved currents and the associated symmetry operators, we first derive the equations of motion. The EoMs are obtained by requiring the action to remain invariant (modulo $2\pi$) under infinitesimal variations of the gauge fields. For instance, under $b^{1}\rightarrow b^{1}+\delta b^{1}$, the condition $\delta S\left[\delta b^{1}\right]=0\mod{2\pi}$ implies $\frac{N_{1}}{2\pi}da^{1}=0$, which is the equation of motion for $a^{1}$. The full set of EoMs, together with the corresponding conserved currents and symmetry operators of $S_{aab}$, is summarized in Table 1. Among these symmetry operators, some are invertible whereas others are non-invertible. In particular, due to the presence of the twisted term, two distinct classes of EoMs and conserved currents emerge, which we refer to as type-I and type-II, respectively, and which will be discussed in detail in the following subsections.
Table 1: Equations of motion (EoMs), conserved currents, and symmetry operators of $S_{aab}=\int\sum_{i=1}^{3}\frac{N_{i}}{2\pi}b^{i}da^{i}+qa^{1}a^{2}b^{3}$, where $q=\frac{pN_{1}N_{2}N_{3}}{\left(2\pi\right)^{2}N_{123}}$ with $p\in\mathbb{Z}_{N_{123}}$. The first three symmetry operators are invertible and correspond to Wilson operators of Abelian particle and Abelian loop excitations. Specifically, $U_{e_{1}}\left(\gamma\right)$, $U_{e_{2}}\left(\gamma\right)$, and $U_{e_{3}}\left(\sigma\right)$ implement $\mathbb{Z}_{N_{1}}^{\left(2\right)}$, $\mathbb{Z}_{N_{2}}^{\left(2\right)}$, and $\mathbb{Z}_{N_{3}}^{\left(1\right)}$ symmetries, respectively, where the superscripts $\left(2\right)$ and $\left(1\right)$ denote 22-form and $1$-form symmetries. In contrast, the last three symmetry operators, $L^{1}\left(\sigma\right)$, $L^{2}\left(\sigma\right)$, and $L^{3}\left(\gamma\right)$, are non-invertible. They take the same functional form as Wilson operators associated with non-Abelian loop and non-Abelian particle excitations, up to overall normalization factors. The delta functions appearing in these symmetry operators enforce constraints on gauge-field configurations, under which the corresponding currents are conserved and define genuine symmetries (see Sec. III.4). The presence of these delta functions renders the symmetry operators non-invertible. Fusion of the symmetry operators listed here generally produces new symmetry operators; see Sec. IV for a detailed discussion.
EoM Currents Continuity Eq. Symmetry operators
@@ III.3 Type-I conserved currents and invertible symmetry operators
We begin by considering a local perturbation of the gauge field $b^{1}$. Under $b^{1}\rightarrow b^{1}+\delta b^{1}$, the variation of the action is $\delta S\left[\delta b^{1}\right]=\int_{M^{4}}\frac{N_{1}}{2\pi}\delta b^{1}da^{1}$. Requiring $\delta S\left[\delta b^{1}\right]=0$ for an arbitrary infinitesimal perturbation $\delta b^{1}$, we obtain the corresponding equation of motion
@@
This EoM implies flux quantization, $\oint_{\gamma}a^{1}=\frac{2\pi k_{1}}{N_{1}}$ with $k_{1}\in\mathbb{Z}_{N_{1}}$; namely, the flux of $a^{1}$ piercing the surface enclosed by $\gamma$ is quantized in integer multiples of $\frac{2\pi}{N_{1}}$. Since this EoM takes the form of a total derivative set to zero, it naturally resembles a continuity equation. Indeed, from Eq. (18) we can extract the continuity equation
@@
which depends only on the topological class of the closed curve $\gamma$. To see that $Q_{1,\gamma}$ is invariant under smooth deformations of $\gamma$, consider $Q_{1,\gamma}-Q_{1,\gamma^{\prime}}=\int_{\gamma-\gamma^{\prime}}*J^{1}_{e}=\int_{\Sigma}d\left(*J^{1}_{e}\right)=0$, where $\Sigma$ is the surface bounded by $\gamma-\gamma^{\prime}$ and we have used Stokes’ theorem. This invariance confirms the topological nature of $Q_{1,\gamma}$.
@@
The object charged under this symmetry is the closed worldsheet of the loop excitation described by $b^{1}$. Correspondingly, $Q_{1,\gamma}$ counts the number of intersections between $\gamma$ and the three-dimensional volume bounded by the loop worldsheet. We therefore identify $U_{e_{1}}\left(\gamma\right)$ as a Wilson operator. Imposing large gauge invariance together with the quantization of $\int_{\gamma}a^{1}$, we find that $e_{1}$ is an integer defined modulo $N_{1}$. As a result, the symmetry operator $U_{e_{1}}\left(\gamma\right)$ with $e_{1}\in\left\{0,1,\cdots,N_{1}-1\right\}$ implements a $\mathbb{Z}_{N_{1}}$ 22-form symmetry.
@@
where $e_{2}\in\left\{1,2,\cdots,N_{2}-1\right\}$ and $e_{3}\in\left\{1,2,\cdots,N_{3}-1\right\}$.

### v3

@@ II Preliminaries of conserved currents and symmetry operators
In relativistic field theory, a continuous symmetry is reflected in a conserved current, and the corresponding conserved charge exponentiates to a topological operator—the symmetry operator. From this viewpoint, a “symmetry” refers to the mathematical structure formed by these topological operators (or, equivalently, by their fusion/commutation relations). In our discussion of generalized symmetries, we adopt the viewpoint that a higher-form symmetry is generated by topological operators originating from higher-form conserved currents. Moreover, this principle naturally extends to the construction of non-invertible symmetries: their symmetry operators can be traced back to a special class of conserved currents (together with additional constraints), as will be discussed in Sec. III. Two representative examples of generalized symmetries in lattice systems are collected in Appendix A.
@@
where the real parameter $\alpha$ takes value in $0,2\pi)$ as a compact phase angle. The compactness of $\alpha$ arises from the quantization of $Q$. The existence of this global symmetry can thus be traced back to the presence of a conserved current. The conservation law ensures that $U_{\alpha}$ is invariant under smooth deformations of $\Sigma_{3}$, hence the terminology of a topological operator for the 0-form (global) symmetry acting on the entire quantum system.
The concept of higher-form symmetry, e.g., $p$-form symmetry, generalizes the above structure by allowing the conserved current to be a $(p+1)$-form. Correspondingly, the associated symmetry operator is supported on a $\left(D-1-p\right)$-dimensional submanifold in (D1+1)(D-1+1)D spacetime. As an illustrative example of higher-form symmetries, we consider the (3+1)(3+1)D Maxwell electromagnetic theory with action $S=\frac{1}{4}\int d^{4}x\,F_{\mu\nu}F^{\mu\nu}$, where $F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$; for simplicity we set the gauge coupling $e=1$ (equivalently, absorb it into the normalization of fields). See also the similar discussion in, e.g., Refs [8, 1, 3, 6]. First, there exists a conserved $2$-form current satisfying $d\left(*J_{m}\right)=0$, with $J_{m}^{\mu\nu}=\frac{1}{4\pi}\epsilon^{\mu\nu\lambda\sigma}F_{\lambda\sigma}=\frac{1}{2\pi}\left(d\widetilde{A}\right)^{\mu\nu}$. Here $\widetilde{A}$ is defined by $d\widetilde{A}=*dA$. This continuity equation is equivalent to $dF=0$, namely the Bianchi identity. The conserved current $J_{m}$ gives rise to a “magnetic” $1$-form symmetry with symmetry operator
@@
Another illustrative example is provided by the mutual Chern–Simons theory, which serves as the effective field theory (i.e., a TQFT description) of a 22d topological order,
@@
The lattice realization of this field theory is the celebrated toric code model in 22d square lattice. Here $a^{1}$ and $a^{2}$ are $1$-form gauge fields. The equations of motion are $\frac{1}{\pi}da^{1}=0$ and $\frac{1}{\pi}da^{2}=0$. In analogy with the previous examples, we interpret these equations as continuity equations and introduce the conserved quantities $Q_{1}=\int_{\gamma}a^{1}$ and $Q_{2}=\int_{\gamma}a^{2}$. These quantities measure the number of topological excitations enclosed by the area bounded by a loop $\gamma$. They generate two minimal symmetry operators (topological line operators),
@@ III Field-theoretical construction of symmetry operators
In this section, we construct symmetry operators (topological operators/defects) from conserved currents in (3+1)(3+1)D TQFTs that serve as effective field-theoretical descriptions of 33d topological orders. We begin with the (3+1)(3+1)D $\mathbb{Z}_{N}$ BFBF theory as a warm-up, which describes $\mathbb{Z}_{N}$ topological orders. We then turn to the main goal of this paper: identifying symmetry operators in the effective field theory of a class of 33d non-Abelian topological orders, namely Borromean-Rings (BR) topological orders. In particular, we show how the conserved-current construction can be exploited to obtain intrinsically non-invertible symmetry operators.
A system with BR topological order [46, 47, 74, 77] exhibits nontrivial BR braiding statistics. A BR braiding process involves one particle and two loop excitations: the particle is moved around the two loops such that its spatial trajectory, together with the two loops, forms Borromean rings. Borromean rings are three closed curves in 33d space that are not pairwise linked, yet cannot be separated as a whole. Consequently, there is no net particle–loop braiding phase, since the particle trajectory is not linked with either loop individually. BR braiding is also distinct from multi-loop braiding because it necessarily involves a particle excitation. The corresponding TQFT is a twisted $BF$ theory with an aabaab-type term; see Eq. (14) below. Remarkably, BR topological order supports non-Abelian particle and loop excitations even when all excitations carry gauge charges and fluxes of an Abelian gauge group, making it an ideal playground for exploring generalized symmetries—including non-invertible ones—in (3+1)(3+1)D topological phases/TQFTs.
In the following, we first derive the equations of motion (EoM) of gauge fields, from which we identify conserved currents. These conserved currents imply generalized symmetries and allow us to construct the corresponding symmetry operators. We find two classes of conserved currents, denoted as type-I and type-II, respectively, as shown in Sec. III.3 and Sec. III.4. Type-I currents lead to group-like (i.e., invertible) higher-form symmetries. By contrast, type-II currents are conserved only under additional constraints such that the corresponding conserved quantities are well-defined. Accordingly, the symmetry operators generated by these conserved quantities must be accompanied by projectors that enforce the constraints. Since projectors have no inverses, these symmetry operators are naturally non-invertible.
@@ III.1 Higher-form symmetries in $\left(3+1\right)$D $\mathbb{Z}_{N}$ BFBF theory
The topological field theories describing the 22d and 33d toric code models discussed above are $\mathbb{Z}_{2}$ BFBF theories. It is straightforward to generalize the gauge group from $\mathbb{Z}_{2}$ to $\mathbb{Z}_{N}$. The (3+1)(3+1)D $\mathbb{Z}_{N}$ BFBF theory is described by
@@
where $J_{e}$ and $J_{m}$ are $3$-form and $2$-form conserved currents, respectively. Analogous to ordinary 0-form symmetries, these conserved currents give rise to conserved quantities,
@@
We now explain the physical meaning of these conserved quantities. Consider $Q_{e}\left(\gamma\right)=\int_{\gamma}a$ as an example. From the correspondence between gauge theory and topological order, the operator $e^{{\rm i}\int_{\gamma}a}$ is the Wilson operator associated with a particle excitation carrying one unit of gauge charge, and $\gamma$ can be interpreted as the worldline of this particle. Acting with $e^{{\rm i}\int_{\gamma}a}$ on a state amounts to transporting the particle along $\gamma$ in spacetime. This process probes the gauge flux carried by loop excitations piercing a surface bounded by $\gamma$.
From the perspective of conserved currents, since $a=a_{\mu}dx^{\mu}$, the current takes the form $J_{e}=-\frac{1}{3!}\epsilon^{\mu\nu\lambda\rho}a_{\mu}\,dx^{\nu}\wedge dx^{\lambda}\wedge dx^{\rho}$. For comparison, in a (3+1)(3+1)D system with a 0-form $U\left(1\right)$ symmetry, the particle current $j=j_{\mu}dx^{\mu}$ is a $1$-form current, whose Hodge dual is $j=j_{0}dx^{1}\wedge dx^{2}\wedge dx^{3}-j_{1}dx^{0}\wedge dx^{2}\wedge dx^{3}+j_{2}dx^{0}\wedge dx^{1}\wedge dx^{3}-j_{3}dx^{0}\wedge dx^{1}\wedge dx^{2}$. Taking $M_{3}$ to be the entire 33d spatial manifold, $\int_{M_{3}}*j$ integrates the particle-number density $j_{0}$ over $M_{3}$, yielding the total particle number. By direct analogy, $\int_{\gamma}*J_{e}$ can be interpreted as a generalized “symmetry charge” associated with the submanifold $\gamma$.
Recall that $\int_{\gamma}*J_{e}$ counts the gauge flux encircled by $\gamma$; in (3+1)(3+1)D, this quantity is given by the linking number between $\gamma$ and the worldsheets $\sigma$ of loop excitations. Equivalently, it can be computed by counting the intersections between $\gamma$ and a three-dimensional volume bounded by $\sigma$ in spacetime. In this sense, $J_{e}$ describes the flow of two-dimensional objects in (3+1)(3+1)D spacetime, in close analogy with $j$ describing the flow of point-like particles. These two-dimensional objects are precisely the worldsheets of loop excitations. The physical interpretation of $Q_{m}\left(\sigma\right)$ follows in a completely analogous manner.
@@
The equations of motion imply that $\oint a\in\frac{2\pi}{N}\mathbb{Z}_{N}$, which in turn gives $NQ_{e}\left(\gamma\right)\in 2\pi\mathbb{Z}$, meaning that $q_{e}$ is equivalent to $q_{e}+N$. Furthermore, large-gauge invariance under $\int_{\gamma}a\rightarrow\int_{\gamma}a+2\pi$ requires $q_{e}$ to be an integer. Consequently, $q_{e}$ takes values in $\mathbb{Z}/N\mathbb{Z}$. The symmetry operators satisfy
@@ III.2 Continuum topological field theory of Borromean-Rings topological order
To identify the conserved currents and the associated symmetry operators, we first derive the equations of motion. The EoMs are obtained by requiring the action to remain invariant (modulo $2\pi$) under infinitesimal variations of the gauge fields. For instance, under $b^{1}\rightarrow b^{1}+\delta b^{1}$, the condition $\delta S\left[\delta b^{1}\right]=0\mod{2\pi}$ implies $\frac{N_{1}}{2\pi}da^{1}=0$, which is the equation of motion for $b^{1}$. The full set of EoMs, together with the corresponding conserved currents and symmetry operators of $S_{\text{BR}}$, is summarized in Table 1. Among these symmetry operators, some are invertible whereas others are non-invertible. In particular, due to the presence of the twisted term, two distinct classes of EoMs and conserved currents emerge, which we refer to as type-I and type-II, respectively, and which will be discussed in detail in the following subsections.
Table 1: Equations of motion (EoMs), conserved currents, and symmetry operators of $S_{\text{BR}}=\int\sum_{i=1}^{3}\frac{N_{i}}{2\pi}b^{i}da^{i}+qa^{1}a^{2}b^{3}$, where $q=\frac{pN_{1}N_{2}N_{3}}{\left(2\pi\right)^{2}N_{123}}$ with $p\in\mathbb{Z}_{N_{123}}$. The first three symmetry operators are invertible and correspond to Wilson operators of Abelian particle and Abelian loop excitations. Specifically, $U_{e_{1}}\left(\gamma\right)$, $U_{e_{2}}\left(\gamma\right)$, and $U_{e_{3}}\left(\sigma\right)$ implement $\mathbb{Z}_{N_{1}}^{\left(2\right)}$, $\mathbb{Z}_{N_{2}}^{\left(2\right)}$, and $\mathbb{Z}_{N_{3}}^{\left(1\right)}$ symmetries, respectively, where the superscripts $\left(2\right)$ and $\left(1\right)$ denote 22-form and $1$-form symmetries. In contrast, the last three symmetry operators, $L^{1}\left(\sigma\right)$, $L^{2}\left(\sigma\right)$, and $L^{3}\left(\gamma\right)$, are non-invertible. They take the same functional form as Wilson operators associated with non-Abelian loop and non-Abelian particle excitations, up to overall normalization factors. The delta functions appearing in these symmetry operators enforce constraints on gauge-field configurations; under these constraints, the corresponding currents are conserved and define genuine symmetries (see Sec. III.4). The presence of these delta functions renders the symmetry operators non-invertible. Fusion of the symmetry operators listed here generally produces new symmetry operators; see Sec. IV for a detailed discussion.
EoMs Currents Continuity Eq. Symmetry operators
@@ III.3 Type-I conserved currents and invertible symmetry operators
We begin by considering a local variation of the gauge field $b^{1}$. Under $b^{1}\rightarrow b^{1}+\delta b^{1}$, the variation of the action is
$\delta S\left[\delta b^{1}\right]=\int_{M^{4}}\frac{N_{1}}{2\pi}\,\delta b^{1}\,da^{1}.$
Requiring $\delta S\left[\delta b^{1}\right]=0$ for an arbitrary infinitesimal perturbation $\delta b^{1}$, we obtain the equation of motion
@@
This EoM implies flux quantization: $\oint_{\gamma}a^{1}=\frac{2\pi k_{1}}{N_{1}}$ with $k_{1}\in\mathbb{Z}_{N_{1}}$. Equivalently, the flux of $a^{1}$ piercing a surface bounded by $\gamma$ is quantized in integer multiples of $\frac{2\pi}{N_{1}}$. Since this EoM is a total-derivative condition, it naturally takes the form of a continuity equation. Indeed, from Eq. (18) we can extract
@@
which depends only on the homology class of the closed curve $\gamma$. To see that $Q_{1,\gamma}$ is invariant under smooth deformations of $\gamma$, consider $Q_{1,\gamma}-Q_{1,\gamma^{\prime}}=\int_{\gamma-\gamma^{\prime}}*J^{1}_{e}=\int_{\Sigma}d\left(*J^{1}_{e}\right)=0$, where $\Sigma$ is a surface with $\partial\Sigma=\gamma-\gamma^{\prime}$ and we have used Stokes’ theorem. This invariance confirms the topological nature of $Q_{1,\gamma}$.
@@
The object charged under this symmetry is the closed worldsheet of the loop excitation described by $b^{1}$. Correspondingly, $Q_{1,\gamma}$ counts the number of intersections between $\gamma$ and a three-dimensional volume bounded by the loop worldsheet. We therefore identify $U_{e_{1}}\left(\gamma\right)$ as a Wilson operator. Imposing large-gauge invariance together with the quantization of $\int_{\gamma}a^{1}$, we find that $e_{1}$ is an integer defined modulo $N_{1}$. As a result, the symmetry operator $U_{e_{1}}\left(\gamma\right)$ with $e_{1}\in\left\{0,1,\cdots,N_{1}-1\right\}$ implements a $\mathbb{Z}_{N_{1}}$ 22-form symmetry.
@@
where $e_{2}\in\left\{0,1,\cdots,N_{2}-1\right\}$ and $e_{3}\in\left\{0,1,\cdots,N_{3}-1\right\}$.